# BRIGHT: A globally distributed multimodal building damage assessment dataset with very-high-resolution for all-weather disaster response

### v2

@@ 2.3 Dataset division strategy
To train ML/DL models using BRIGHT and evaluate their generalizability, it is necessary to divide the dataset into a training set, validation set, and test set. Gerard et al. (2024) suggested that dividing the dataset on an event-by-event basis, rather than randomly across the entire dataset, provides a more accurate reflection of a model’s generalizability. Therefore, for the 14 events listed in Table 2, we divide the corresponding data for each event into training, validation, and test sets in a ratio of 7:1:2. The sets obtained for each event are then merged to create the final training, validation, and test sets. In our experiments, DL models are trained using the training set, and the optimal hyperparameters (e.g., learning rate) are selected based on performance on the validation set. The generalization capability of the models is subsequently evaluated on the test set.
In addition to the above standard ML division, we also introduce a cross-event transfer setup to better evaluate the ability of models to generalize across disaster events, a critical challenge in real-world applications where models are expected to handle unseen disaster types and locations. This setup comprises two key divisions:
@@
One-shot setup: Recognizing the difficulty of the zero-shot setup, we introduce a one-shot setup, acknowledging that in real-world disaster response, a small number of labeled samples from the new disaster can be obtained. In this setting, a limited subset of labeled data (one image for training and one image for validation) from the target disaster event is incorporated into the training process. At the same time, the majority of the test set remains unseen. This setup evaluates the model’s ability to leverage a minimal amount of manually labeled data to improve disaster-specific adaptation.
@@ 3.1 Problem statement and two solutions
The objective of building damage assessment is to interpret EO data covering the affected area and generate a building damage proxy map that reflects the extent of damage to buildings. To achieve this, two common solutions are typically employed.
One is to directly treat the building damage assessment task as a single semantic segmentation task (Adriano et al., 2021; Gupta and Shah, 2021). The pre- and post-event images are taken as inputs of the model, and then the final damage map is directly predicted, which can be formalized as , where is is the pre-event imagery, is the post-event imagery, is a semantic segmentation model, is the obtained damage map. In the context of this paper, is VHR optical imagery and is VHR SAR imagery.
@@
Since the aim of this article is not only to provide a large-scale multimodal benchmark dataset to support all-weather disaster response but also to offer insights for designing appropriate methods in future research, we will employ and compare both solutions in our experiments.
@@ 3.2 Benchmark suites
We evaluate BRIGHT using several advanced deep network architectures from both the computer vision and remote sensing communities. Since the building damage assessment task can be considered a specialized semantic segmentation task, we selected two well-known segmentation networks from the computer vision field: UNet (Ronneberger et al., 2015) and DeepLabV3+ (Chen et al., 2018). Additionally, we included several state-of-the-art networks from the EO community: SiamAttnUNet (Adriano et al., 2021), SiamCRNN (Chen et al., 2020), ChangeOS (Zheng et al., 2021), DamageFormer (Chen et al., 2022a), and ChangeMamba (Chen et al., 2024). These networks encompass a broad range of representative deep learning paradigms, including CNNs, RNNs, Transformers, and the more recent Mamba architecture.
Among these seven networks, UNet, DeepLabV3+, and SiamAttnUNet adopt the first solution, directly treating building damage assessment as a single semantic segmentation task. In contrast, SiamCRNN, ChangeOS, DamageFormer, and ChangeMamba adopt the second solution by decoupling the task into two subtasks: building localization and damage classification.
@@ 3.3 Model training
For UNet, DeepLabV3+, and SiamAttnUNet, which directly predict damage proxy maps from the input multimodal image pairs, the training loss function is defined as:
@@
For zero-shot and one-shot setups, we incorporate UDA and SSL techniques to better utilize limited or unlabeled target disaster data. The specifics of these training adjustments, along with the UMCD and UMIM approaches, are provided in Appendix E, F, and G, respectively.
@@ 3.4 Accuracy assessment
We adopt overall accuracy (OA), F1 score (F1), and mean intersection over union (mIoU) to evaluate the performance of DL models on building damage assessment. These metrics are commonly used in previous research on building damage assessment (Zheng et al., 2021). Following the setup in prior unimodal building damage assessment studies and the related xView2 Challenge (Gupta et al., 2019), F1 score is used to assess the performance of DL models in the building localization and damage classification subtasks. Finally, OA and mIoU are employed to measure the overall quality of the proxy damage map, providing a comprehensive assessment of the models’ ability to localize buildings and classify damage levels accurately.
@@ 4.1 Evaluation results on standard machine learning split
Table 5 shows the evaluation results for each model on the test set. We observe that ChangeMamba achieves the best overall performance, with an OA of 96.22, an mIoU of 67.63, and the highest and scores (90.90 and 72.70, respectively). DamageFormer also performs well, closely following ChangeMamba, with an mIoU of 67.09 and an OA of 96.13. Both models demonstrate strong capability in both building localization and damage classification tasks. The accuracy advantage of ChangeMamba and DamageFormer underscores the importance of leveraging advanced deep learning architectures to improve performance in complex tasks such as building damage assessment. For models using the direct prediction approach (UNet, DeepLabV3+, SiamAttnUNet), UNet delivers the best results, with an mIoU of 64.94 and an OA of 95.47. However, its performance still lags behind that of the decoupled models, further emphasizing the advantage of task decoupling.
@@
Finally, Figure 6 shows some of the building damage proxy maps obtained by the seven models on the test set.
@@ 4.2 What have models learned and what can they learn?
To better understand model behavior beyond performance metrics, we explore the internal attention patterns of trained models using class activation maps (CAMs). Figure 7 presents the CAM responses of ChangeOS across six representative disaster events, including volcano, explosion, flood, wildfire, earthquake, and hurricane. We observe that the attention distribution varies across disaster types. In Goma-VE-2021, for the “Destroyed” category, the encoder processing the optical imagery exhibits strong activations across nearly all built-up regions, accurately localizing individual buildings. This suggests the model has effectively learned to extract detailed structural cues from pre-event optical imagery. In contrast, the encoder handling the SAR imagery shows intense activation over the lava-covered region on the left, indicating that the model has identified this as a key signal for destruction—likely due to the significant backscatter changes caused by lava flow. In the “Damaged” category, however, the activations are more subtle. Attention is primarily focused near the boundary of the lava flow, where partial or ambiguous structural changes occur. In the decoder stage, the destroyed buildings in the lava-affected area are strongly activated, aligning well with the reference labels. Conversely, for the “Damaged” class, only a few regions are activated. This suggests that inferring partial damage from SAR imagery in volcanic disaster scenarios remains a significant challenge, as the subtle structural degradation is often not clearly reflected in SAR backscatter or texture.
@@
Beyond understanding what the models have already learned, a more critical question is: What can they learn? More specifically, to what extent can optical–SAR combinations support accurate building damage assessment across diverse disaster scenarios? To investigate this, we aggregate the mIoU of the seven models and show their average mIoU across seven major disaster types in Figure 8. Most notably, wildfire and volcano events exhibit the highest IoU scores for the “Destroyed” category, both exceeding 70. This indicates that the model can effectively leverage SAR-based backscatter anomalies, such as lava flows or widespread debris fields, to detect fully destroyed structures. These results demonstrate the strong potential of Optical–SAR fusion in such high-impact scenarios. However, for the “Damaged” category, the performance drops significantly. The average IoU for damaged buildings in wildfire events remains below 20, while in volcano events, it falls to around 5. This suggests that single-polarization SAR imagery lacks the fine-grained information needed to reliably distinguish partially damaged buildings, where structural integrity may still be partially preserved and backscatter signals remain ambiguous. For earthquake events, both damaged and destroyed categories yield relatively low IoUs. This is likely due to the complex and heterogeneous patterns of structural collapse typical of seismic events, where damage is often subtle, partial, and highly variable. These conditions pose significant challenges for SAR-based assessment. Interestingly, for flood and hurricane events, the model achieves relatively high IoU scores for the “Damaged” category with about 50 and 60, highlighting that SAR is effective at capturing contextual environmental changes—such as water inundation or terrain disruption—which indirectly aid in assessing building damage. In the case of the conflict event, the model’s performance on the destroyed class is surprisingly low. We attribute this to the limited number of destroyed samples in the dataset for this category, which leads to insufficient learning and poor generalization.
@@ 4.3 Effect of post-processing methods
Post-processing techniques help refine the raw predictions from deep learning models, reducing noise, improving consistency, and ensuring spatial coherence in damage proxy maps (Zheng et al., 2021). Here, we further explore the impact of post-processing algorithms. Table 7 presents the effect of three post-processing techniques applied to ChangeMamba. The post-processing methods evaluated include test-time augmentation, object-based major voting, and model ensembling. Further details on these methods are provided in Appendix D
@@
4.4 Evaluation results on cross-event transfer split
4.4.1 Baseline methods
@@
Cross-event transfer, especially under zero-shot settings, poses a significant challenge for building damage assessment. As shown in Table 8, the average mIoU across all baseline models in the zero-shot setting is below 40, a noticeable drop compared to their performance under the fully supervised standard ML split, where models typically achieve 48 to 52 mIoU. This performance gap underscores the difficulty of generalizing to unseen disaster events without access to any target-domain supervision due to substantial domain shifts in imaging conditions, damage patterns, urban morphology, and sensor response. Despite this, all models exhibit clear performance gains in the one-shot setting, where a small number of labeled samples from the target event are available. This suggests that even minimal supervision can significantly aid adaptation to new disaster contexts.
@@ 4.4.2 Unsupervised domain adaptation and semi-supervised learning methods
Table 9 reports the performance of five classical UDA methods on the BRIGHT dataset. Compared to the source-only baseline (35.60 mIoU), most methods achieve modest improvements, with FDA (+0.69) and AdaptSeg (+0.45) showing the most consistent gains. CLAN also performs slightly better (+0.41), suggesting that category-level alignment contributes positively even under a large domain shift. In contrast, AdvEnt shows negligible change, and PyCDA significantly underperforms, dropping over 2 mIoU compared to the baseline. This performance degradation indicates that approaches relying on pseudo-label refinement or curriculum learning may struggle in disaster imagery, where spatial layout and damage semantics vary drastically across events. Overall, while UDA methods show some promise, their improvements are relatively minor and not robust across all events, underscoring the difficulty of domain alignment in EO-based damage mapping.
In the one-shot setting, we evaluate several popular SSL methods, summarized in Table 10. With only minimal supervision in the target event, all methods except CPS achieve consistent improvements over the supervised one-shot baseline (38.82). The best-performing method is Mean Teacher (MT), which yields a gain of +1.18, followed by CCT (+0.67) and GCT (+0.52). These results show that simple consistency-based teacher-student frameworks are particularly effective in leveraging unlabeled data under limited supervision, likely due to their robustness to noisy or class-imbalanced targets.
4.4.3 Why cross-event transfer is challenging?
@@
Inconsistent damage signatures across events. Figure 9 presents violin plots of SAR backscatter values for intact, damaged, and destroyed buildings across multiple event pairs. These plots reveal two key observations. Firstly, even within the same disaster category (the first row), the pixel intensity distributions for damaged and destroyed buildings differ significantly between events. This indicates that SAR-based damage signatures are not consistent across locations, possibly due to differences in urban layout, building materials, or sensor incidence angles. In addition, the distribution shift becomes more pronounced across different disaster types (the second row). For instance, the destroyed-building signatures in wildfires are notably different from those in volcano eruptions, which again differ from floods or hurricanes. These variations reflect fundamental differences in damage mechanisms: buildings burned in wildfires, submerged in floods, or collapsed in earthquakes leave very different patterns in SAR backscatter. Such distributional discrepancies make it extremely difficult for models to generalize from one event to another. A model trained on one disaster may learn class boundaries (e.g., between “Damaged” and “Destroyed”) that do not transfer well to another disaster, especially when the visual and physical properties of damage are fundamentally different.
@@
Lack of target domain supervision for model selection. Another critical challenge in cross-event setup is the absence of labeled target-domain samples for model selection. In typical domain adaptation benchmarks in computer vision, a target domain validation set is available to tune hyperparameters or select the best checkpoint. However, in disaster response scenarios, no such validation data can be assumed. Figure 10 illustrates the resulting issue: the gap between the actual best mIoU on the test event and the mIoU of the checkpoint selected using a source-domain validation set for several representative models under both zero-shot and one-shot settings. In the zero-shot setting, this performance gap is substantial across most models. This indicates that relying on source-domain validation leads to suboptimal model selection due to poor reflection of the target distribution. The one-shot setting helps reduce this gap by enabling limited target-aware selection. However, the issue is not entirely resolved. Even with a few labeled samples, model instability and domain shift still make selection challenging.
Together, these two findings highlight that cross-event transfer is hindered not only by semantic and statistical shifts across disaster types but also by operational constraints that prevent ideal training and tuning. Future solutions should consider both aspects—developing models that are robust to distributional variance and designing selection strategies that do not depend on target supervision, such as self-validation, early stopping heuristics, or domain-agnostic performance proxies (Yang et al., 2023).

### v3

@@ 2.3 Dataset division strategy
To train DL models using Bright and evaluate their generalizability, it is necessary to divide the dataset into a training set, validation set, and test set. Gerard et al. (2024) suggested that dividing the dataset on an event-by-event basis, rather than randomly across the entire dataset, provides a more accurate reflection of a model’s generalizability. Therefore, for the 14 events listed in Table 2, we divide the corresponding data for each event into training, validation, and test sets in a ratio of 7:1:2. The sets obtained for each event are then merged to create the final training, validation, and test sets. In our experiments, DL models are trained using the training set, and the optimal hyperparameters (e.g., learning rate) and checkpoints are selected based on performance on the validation set. The generalization capability of the models is subsequently evaluated on the test set.
In addition to the above standard machine learning division, we also introduce a cross-event transfer setup to better evaluate the ability of models to generalize across disaster events, a critical challenge in real-world applications where models are expected to handle unseen disaster types and locations. This setup comprises two key divisions:
@@
One-shot setup: Recognizing the difficulty of the zero-shot setup, we introduce a one-shot setup, acknowledging that in real-world disaster response, a small number of labeled samples from the new disaster can be obtained. In this setting, a limited subset of labeled data (one pair for training and one pair for validation) from the target disaster event is incorporated into the training process. At the same time, the majority of the test set remains unseen. This setup evaluates the model’s ability to leverage a minimal amount of manually labeled data to improve disaster-specific adaptation.
@@ 3.1 Problem statement and two solutions
The objective of building damage assessment is to interpret EO data covering the affected area and generate a building damage map that reflects the extent of damage to buildings. To achieve this, two common solutions are typically employed.
One is to directly treat the building damage assessment task as a single semantic segmentation task (Adriano et al., 2021; Gupta and Shah, 2021). The pre- and post-event images are taken as inputs of the model, and then the final damage map is directly predicted. This process can be formalized as , where is is the pre-event imagery, is the post-event imagery, is a semantic segmentation model, is the obtained damage map. In the context of this paper, is VHR optical imagery and is VHR SAR imagery.
@@
Since the aim of this article is not only to provide a large-scale multimodal dataset to support all-weather disaster response but also to offer insights for designing appropriate methods in future research, we will employ and compare both solutions in our experiments.
@@ 3.2 Benchmark suites
We evaluate several advanced deep network architectures from both the computer vision and EO communities on Bright. Since the building damage assessment task can be considered a specialized semantic segmentation task, we selected two well-known segmentation networks from the computer vision field: UNet (Ronneberger et al., 2015) and DeepLabV3+ (Chen et al., 2018). Additionally, we included several state-of-the-art networks from the EO community: SiamAttnUNet (Adriano et al., 2021), SiamCRNN (Chen et al., 2020), ChangeOS (Zheng et al., 2021), DamageFormer (Chen et al., 2022a), and ChangeMamba (Chen et al., 2024). These networks encompass a broad range of representative DL architectures, including convolutional neural networks (CNNs), recurrent neural networks (RNNs), Transformers, and the more recent Mamba architecture.
Among these seven networks, UNet, DeepLabV3+, and SiamAttnUNet adopt the first solution, directly treating building damage assessment as a single semantic segmentation task. In contrast, SiamCRNN, ChangeOS, DamageFormer, and ChangeMamba adopt the second solution by decoupling the task into building localization and damage classification tasks.
@@ 3.3 Model training
For UNet, DeepLabV3+, and SiamAttnUNet, which directly predict damage maps from the input multimodal image pairs, the training loss function is defined as:
@@
For zero-shot and one-shot setups, we incorporate UDA and SSL techniques to better utilize unlabeled and limited target disaster data. The specifics of these training adjustments, along with the UMCD and UMIM approaches, are provided in Appendix E, F, G, and H, respectively.
@@ 3.4 Accuracy assessment
We adopt overall accuracy (OA), F1 score (F1), and mean intersection over union (mIoU) to evaluate the performance of DL models on building damage assessment. These metrics are commonly used in previous research on building damage assessment (Zheng et al., 2021). Following the setup in prior unimodal building damage assessment studies and the related xView2 Challenge (Gupta et al., 2019), the F1 score is used to assess the performance of DL models in the building localization and damage classification subtasks. Finally, OA and mIoU are employed to measure the overall quality of the building damage map, providing a comprehensive assessment of the models’ ability to localize buildings and classify damage levels accurately.
@@ 4.1 Evaluation results on standard machine learning split
Table 5 shows the evaluation results for each model on the test set. We observe that ChangeMamba achieves the best overall performance, with an OA of 96.22, an mIoU of 67.63, and the highest and scores (90.90 and 72.70, respectively). DamageFormer also performs well, closely following ChangeMamba, with an mIoU of 67.09 and an OA of 96.13. Both models demonstrate strong capability in both building localization and damage classification tasks. The accuracy advantage of ChangeMamba and DamageFormer underscores the importance of leveraging advanced DL architectures to improve performance in complex tasks such as building damage assessment. For models using the direct prediction approach (UNet, DeepLabV3+, SiamAttnUNet), UNet delivers the best results, with an mIoU of 64.94 and an OA of 95.47. However, its performance still lags behind that of the decoupled models, further emphasizing the advantage of task decoupling.
@@
Finally, Figure 6 shows some of the building damage maps obtained by the seven models on the test set.
@@ 4.2 What have models learned and what can they learn?
To better understand model behavior beyond performance metrics, we explore the internal attention patterns of trained models using class activation maps (CAMs). Figure 7 presents the CAM responses of ChangeOS across six representative disaster events, including volcano, explosion, flood, wildfire, earthquake, and hurricane. We observe that the attention distribution varies across disaster types. Taking Goma-VE-2021 as the example, for the “Destroyed” category, the encoder processing the optical imagery exhibits strong activations across nearly all built-up regions, accurately localizing individual buildings. This suggests the model has effectively learned to extract detailed structural cues from pre-event optical imagery. In contrast, the encoder handling the SAR imagery shows intense activation over the lava-covered region on the left, indicating that the model has identified this as a key signal for destruction—likely due to the significant backscatter changes caused by lava flow. In the “Damaged” category, however, the activations are more subtle. Attention is primarily focused near the boundary of the lava flow, where partial or ambiguous structural changes occur. In the decoder stage, the destroyed buildings in the lava-affected area are strongly activated, aligning well with the reference labels. Conversely, for the “Damaged” class, only a few regions are activated. This suggests that inferring partial damage from SAR imagery in volcanic disaster scenarios remains a significant challenge, as the subtle structural degradation is often not clearly reflected in SAR backscatter or texture.
@@
Beyond understanding what the models have already learned, a more critical question is: What can they learn? More specifically, to what extent can optical–SAR combinations support accurate building damage assessment across diverse disaster scenarios? To investigate this, we aggregate the mIoU of the seven models and show their average mIoU across seven major disaster types in Figure 8. Most notably, wildfire and volcano events exhibit the highest IoU scores for the “Destroyed” category, both exceeding 70. This indicates that the model can effectively leverage SAR-based backscatter anomalies, such as lava flows or widespread debris fields, to detect fully destroyed structures. These results demonstrate the strong potential of optical–SAR fusion in such high-impact scenarios. However, for the “Damaged” category, the performance drops significantly. The average IoU for damaged buildings in wildfire events remains below 20, while in volcano events, it falls to around 5. This suggests that single-polarization SAR imagery lacks the fine-grained information needed to reliably distinguish partially damaged buildings, where structural integrity may still be partially preserved and backscatter signals remain ambiguous. For earthquake events, both damaged and destroyed categories yield relatively low IoUs. This is likely due to the complex and heterogeneous patterns of structural collapse typical of seismic events, where damage is often subtle, partial, and highly variable. These conditions pose significant challenges for SAR-based assessment. Interestingly, for flood and hurricane events, the model achieves relatively high IoU scores for the “Damaged” category with about 50 and 60, highlighting that SAR is effective at capturing contextual environmental changes, such as water inundation or terrain disruption, which indirectly aid in assessing building damage. In the case of the conflict event, the model’s performance on the “Destroyed” class is surprisingly low. We attribute this to the limited number of destroyed samples in the dataset for this category, which leads to insufficient learning and poor generalization.
@@ 4.3 Effect of post-processing methods
Post-processing techniques help refine the raw predictions from DL models, reducing noise, improving consistency, and ensuring spatial coherence in damage maps (Zheng et al., 2021). Here, we further explore the impact of post-processing algorithms. Table 7 presents the effect of three post-processing techniques applied to ChangeMamba. The post-processing methods evaluated include test-time augmentation, object-based major voting, and model ensembling. Further details on these methods are provided in Appendix D.
@@
4.4 Evaluation results on cross-event transfer setup
4.4.1 Baseline methods
Cross-event transfer, especially under zero-shot settings, poses a significant challenge for building damage assessment. As shown in Table 8, the average mIoU across all baseline models in the zero-shot setting is below 40, a noticeable drop compared to their performance under the fully supervised standard ML split, where models typically achieve 48 to 52 mIoU. This performance gap underscores the difficulty of generalizing to unseen disaster events without access to any target domain supervision due to substantial domain shifts in imaging conditions, damage patterns, urban morphology, and sensor response. Despite this, all models exhibit clear performance gains in the one-shot setting, where a small number of labeled samples from the target event are available. This suggests that even minimal supervision can significantly aid adaptation to new disaster contexts.
@@ 4.4.2 Unsupervised domain adaptation and semi-supervised learning methods
Table 9 reports the performance of five classical UDA methods. Compared to the source-only baseline (35.60 mIoU), most methods achieve modest improvements, with FDA (+0.69) and AdaptSeg (+0.45) showing the most consistent gains. CLAN also performs slightly better (+0.41), suggesting that category-level alignment contributes positively even under a large domain shift. In contrast, AdvEnt shows negligible change, and PyCDA significantly underperforms, dropping over 2 mIoU compared to the baseline. This performance degradation indicates that approaches relying on pseudo-label refinement or curriculum learning may struggle in multimodal imagery under disaster scenarios, where spatial layout and damage semantics vary drastically across events. Overall, while UDA methods show some promise, their improvements are relatively minor and not robust across all events, underscoring the difficulty of domain alignment in EO-based damage mapping.
In the one-shot setting, we evaluate several popular SSL methods, summarized in Table 10. With only minimal supervision in the target event, all methods except CPS achieve improvements over the one-shot baseline (38.82). The best-performing method is Mean Teacher (MT), which yields a gain of +1.18, followed by CCT (+0.67) and GCT (+0.52). These results show that simple consistency-based teacher-student frameworks are particularly effective in leveraging unlabeled data under limited supervision, likely due to their robustness to noisy or class-imbalanced targets.
4.4.3 Why is cross-event transfer challenging?
@@
Inconsistent damage signatures across events. Figure 9 presents violin plots of SAR backscatter values for intact, damaged, and destroyed buildings across multiple event pairs. These plots reveal two key observations. Firstly, even within the same disaster category (the first row), the pixel intensity distributions for damaged and destroyed buildings differ significantly between events. This indicates that SAR-based damage signatures are not consistent across locations, possibly due to differences in urban layout, building materials, or sensor incidence angles. Secondly, the distribution shift becomes more pronounced across different disaster types (the second row). For instance, the signatures of destroyed building in wildfires are notably different from those in volcano eruptions, which again differ from floods or hurricanes. These variations reflect fundamental differences in damage mechanisms: buildings burned in wildfires, submerged in floods, or collapsed in earthquakes leave very different patterns in SAR backscatter. Such distributional discrepancies make it extremely difficult for models to generalize from one event to another. A model trained on one disaster may learn class boundaries (e.g., between “Damaged” and “Destroyed”) that do not transfer well to another disaster, especially when the visual and physical properties of damage are fundamentally different.
@@
Lack of target domain supervision for model selection. Another critical challenge in cross-event setup is the absence of labeled target domain samples for model selection. In typical domain adaptation benchmarks in computer vision, a target domain validation set is available to tune hyperparameters or select the best checkpoint. However, in disaster response scenarios, no such validation data can be assumed. Figure 10 illustrates the resulting issue: the gap between the actual best mIoU on the test event and the mIoU of the checkpoint selected using a source domain validation set for several representative models under both zero-shot and one-shot settings. In the zero-shot setting, this performance gap is substantial across most models. This indicates that relying on source domain validation leads to suboptimal model selection due to poor reflection of the target distribution. The one-shot setting helps reduce this gap by enabling limited target-aware selection. However, the issue is not entirely resolved. Even with a few labeled samples, model instability and domain shift still make selection challenging.
Together, these two findings highlight that cross-event transfer is hindered not only by semantic and statistical shifts across disaster types but also by operational constraints that prevent ideal training and tuning. Future solutions should consider both aspects, i.e., developing models that are robust to distributional variance and designing selection strategies that do not depend on target supervision, such as self-validation, early stopping heuristics, or domain-agnostic performance proxies (Yang et al., 2023).